OpenStudy (anonymous):

use reference angles to find exact value of the expression sin -14 pi over 3

OpenStudy (tkhunny):

What is the "Reference Angle"? -14pi/3 + 6pi/3 = -8pi/3 -8pi/3 + 6pi/3 = -2pi/3 -2pi/3 + 6pi/3 = 4pi/3 Two questions: 1) What in the World am I doing with 2pi? Why is that useful? 2) What is the reference angle?